OpenStudy (anonymous):

Calculate the mean and standard deviation of the probability distribution created by rolling (6 sided)die.

Directrix (directrix):

If a fair die is thrown, there are six possible outcomes from 1 to 6, and each is equally likely. The mean is 3.5, the variance is 35/12, and the standard deviation is 1.71. The mean =(1 + 6) / 2 = 7/2 = 3.5 where 1 and 6 are the lower extreme and upper extremes of the distribution.

OpenStudy (amistre64):

i was thinking it was a uniform distribution

Directrix (directrix):

Calculate the standard deviation. Use the formula Sqrt[ ( n^2 - 1)/12] where n = 6. Sqrt((36-1)/12)) = Sqrt [ (35)/12] = 1.71.